# Rethinking Detection Based Table Structure Recognition for Visually Rich Documents

### 01 Dec 2023 v1

@@ 3.4 Spatial Attention and Deformable Convolution
As discussed in Section 1, building long dependencies for detecting the defined components is important. Inspired by the recent studies using large convolution kernels ding2022scaling ; liu2022convnet , a Spatial Attention Module has been proposed, as shown in Figure 6. For the design of the Spatial Attention Module, we use a similar architecture with MSCA guo2022segnext containing multiple branches and large kernel convolutions and use spatial and depthwise separable convolution chollet2017xception ; howard2017mobilenets to reduce the number of parameters. More specifically, for the spatial separable convolution, we use a pair of $7*1$ and $1*7$ kernels to replace a typical $7*7$, use the pair of $11*1$ and $1*1$ 11, and the pair of $21\times 1$ and $1\times 21$ to replace $11*1$ 11 and $21*21$ kernels, respectively. For the depthwise separable convolution, we applied the convolution on each channel of the feature maps independently. Then, the outputs of the three branches are concatenated together as the input of a convolution layer with $1\times 1$ kernel to make the channel dimension the same as the inputs. The proposed Spatial Attention Module can be easily inserted into the backbone network between two blocks because they do not change the feature shapes. For example, for a typical backbone network implemented by ResNet he2016deep containing a STEM block and four Residual Blocks, as shown in Figure 6, the Spatial Attention module can be inserted after the last three Residual Blocks to generate the spatial attention, then the spatial attention can be applied to the original outputs of each Residual Block by Element-wise Multiplication. It is worth mentioning that the Spatial Attention Modules shown in Figure 6 have independent trainable parameters, and all the feature maps are padded correspondingly to keep the size of the feature maps.
@@
On the other hand, deformable convolution dai2017deformable has been demonstrated to help improve the detection performance on the document image detection tasks by many studies fernandes2023tablestrrec ; siddiqui2018decnt ; mondal2023dataset . Deformable convolution uses a learnable grid offset to sample the grid points from the feature map and apply the convolution operation, as defined by Equation 6,
$\mathbf{z}_{p_{0}}=\sum_{p_{n}\in R}{w(p_{n})x(p_{0}+p_{n}+\Delta p_{n})}$ (6)
where $p_{0}$ is the location on the output feature map $\mathbf{z}$, $p_{n}$ is the $n$th grid point in grid $R$, and $\Delta p_{n}$ is the $n$th learnable offset. Since the offset $\Delta p_{n}$ applied to the deformable convolution is usually obtained by a regular convolution with small kernels, such as a 3 * 3 kernel, it can only improve the local feature instead of building long dependencies. Therefore, deformable convolution is useful for improving the detection performance but can contribute little to exploring the structures of the table. In this study, we apply the proposed Spatial Attention module and deformable convolution to build the long-dependent and local features together.
@@ 4.1 Datasets and Experimental Settings
We utilize four datasets in this work, including SciTSR chi2019complicated , FinTabNet zheng2020global , PubTabNet zhong2020image and PubTables1M smock2022pubtables . As discussed in the study in smock2023aligning , FinTabNet and SciTSR datasets contain noise annotations that are harmful for the model performance. Therefore, we use their cleaned versions proposed in the study in smock2023aligning . All the four datasets consist of only the table regions of each image. SicTSR dataset is collected from academic publications, which contains 7453, 1034 and 2134 samples for the training, validation and testing. PubTables1M dataset is a large scale dataset for the TSR problem collected from the PMCOA corpus, containing 758849 training samples, 94959 validation samples, and 93834 testing samples. Since the PubTabNet dataset does not provide original PDF files, we cannot process it to make detection annotations. Besides, its testing is not public available. Therefore, we use its validation set to evaluate the model trained with the PubTable1M dataset. Following the study in smock2023aligning , we use the code base in smock2021tabletransformer to process the datasets and align the formats of all the three datasets. FinTabNet is also a large dataset widely used for the TSR problem, containing 78537, 9289 and 9650 samples for the training, testing and validation, respectively. FinTabNet is collected from the annual reports of companies, making its data source different from other two datasets. Table 1 summarizes the datasets used in this study for the model evaluation.
@@
Since the TSR problem in this study is formulated as an object detection problem, we use both detection metrics and cell-level TSR metric for the model evaluation. For the detection metrics, we employ the widely accepted COCO metrics lin2014microsoft , namely mean Average Precision (mAP), $AP50$, $AP75$, $AP_{s}$, $AP_{m}$, $AP_{l}$, and object specific AP scores. More specifically, $AP50$, $AP75$ are the APs using 0.50 and 0.75 as IoU thresholds, respectively. $AP_{s}$, $AP_{m}$ and $AP_{l}$ are the APs of different size of target objects, which is defined by Equation 7.
@@
For the TSR metric, we choose structure-only Tree-Edit-Distance-Based Similarity(TEDS) zhong2020image , which is firstly introduced in the study to overcome the drawbacks of adjacency relation metrics hurst2003constraint , and can be defined as Equation 8.
where EditDist is the tree-edit distance, and $T$ is the number of nodes in the tree. It is worth mentioning that we use structure-only TEDS in the study, which only consider the HTML tags without extracting their contents to avoid the influence of OCR tools. It is worth mentioning that the testing samples can be categorized into Sim. and Com. groups based on whether they have cells spanning across multiple columns and rows.
@@ 4.2 Implementation Details and Experimental Results
To verify the effectiveness of our proposed solution, we include all three types of state-of-the-art methods as discussed in Section 2. For the detection-based methods, Cascade R-CNN Cai2018CascadeRD , Deformable-DETR zhu2020deformable and Sparse R-CNN sun2021sparse as benchmark models, in which Cascade R-CNN Cai2018CascadeRD is also the based model of the proposed methods, Deformable-DETR and Sparse R-CNN are two state-of-the-art transformer based models. For the image-to-sequence models, EDD zhong2020image , GTE zheng2020global , TableFormer nassar2022tableformer , TableMaster ye2021pingan , VAST huang2023improving and MTL-TabNet visapp23namly are included. TSRFormer-DQ-DETR WANG2023109817 is a state-of-the-art model following the pipeline of detecting separation lines and then merging cell grids, as discussed in Section 2.
We implement Cascade R-CNN and our proposed TSRDet based on the Detection2 wu2019detectron2 , build the Deformable-DETR based on detrex ren2023detrex , train the Sparse R-CNN with their official codebase. For the Deformable-DETR and Sparse R-CNN, we use the default parameters. For the Cascade R-CNN baseline, we align the number of regional proposals and the batch normalization method to the TSRDet, as shown in Table 2. All these detection models are using ResNet50 he2016deep pre-trained with ImageNet deng2009imagenet as the backbone network. We also re-train TableMaster ye2021pingan with the FinTabNet dataset based on its official code base. For the implementation of the proposed TSRDet, aspect ratios in the anchor box generation are set as [0.0125, 0.025, 0.0625, 0.125, 0.25, 0.5, 1.0, 2.0, 4.0, 8.0, 16, 40, 80], and other key parameters are summarized in Table 2. Notably, to calculate the structure-only TEDS, we use the scripts provided by study in smock2021tabletransformer to generate the HTML sequences from the detected components. All the models are trained with 240, 120, and 60 epochs for the SciTSR, FinTabNet, and PubTables1M datasets, respectively.
The experimental results regarding the COCO metrics and structure-only TEDS are shown in Tables 7,  34,  5 and  6, which can demonstrate the superiority of the proposed solution. For the SciTSR dataset, the proposed TSRDet can improve the baseline Cascade R-CNN by 19.32% regarding the structure-only TEDS, outperforming Deformable-DETR and Sparse R-CNN. When it comes to COCO metrics, the mAP of the proposed TSR is as good as Deformable-DETR, outperforming other benchmark models. Similarly, the proposed TSR can outperform benchmark models regarding both COCO metrics and structure-only TEDS on the FinTabNet and PubTables1M datasets. It is worth mentioning that FinTabNet does not provide original PDF files, making it hard to generate detection annotations. Therefore, the model performance reported in Table 6 is calculated using the model trained with the PubTable1M dataset. Even though the PubTable1M and PubTabNet datasets have misalignments regarding the ground truth HTML sequences, the proposed method still shows competitive performance compared with other state-of-the-art methods, as shown in Table 6. Figure 7 shows a prediction sample and its generated HTML sequence after post-processing, demonstrating the capacities of the proposed solution.
Besides demonstrating the effectiveness of the proposed TSRDet, there are some important observations from the experimental results. First, structure-only TEDS is not well aligned with COCO metrics. For example, even though the baseline Cascade R-CNN model can only reach 87.49% regarding the structure-only TEDS on the FinTabNet dataset, its detection performance is only slightly worse than other models. Deformmable-DETR has better mAP, but worse structure-only TEDS than Sparse R-CNN model on the FinTabNet dataset. Second, benchmark models, especially Sparse R-CNN also have very promising performance even though it is not as good as our proposed solution. The discussions and analysis to these observations will be in Section 4.4.
@@
PRE_NMS_TOPK_TRAIN 4000 RPN proposals to keep before applying NMS
PRE_NMS_TOPK_TEST 2000 RPN proposals to keep before applying NMS
POST_NMS_TOPK_TRAIN 4000 RPN proposals to keep after applying NMS
POST_NMS_TOPK_TEST 2000 RPN proposals to keep after applying NMS
@@
Dataset Model $AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$ table column row spanning cell projected row header column header SciTSR Cascade R-CNN 93.89 95.27 94.80 95.81 93.89 92.96 98.96 98.63 96.33 88.58 83.80 97.01 Deformable-DETR 96.28 97.39 97.01 96.75 96.55 96.07 98.96 98.63 97.26 93.84 90.86 98.15 Sparse-RCNN 94.78 96.17 95.48 95.49 95.07 90.08 98.98 98.30 97.93 88.06 86.92 98.49 TSRDet(Ours) 96.28 96.79 96.57 99.01 96.42 95.65 98.97 99.25 98.57 95.30 87.06 98.50 FinTabNet Cascade R-CNN 95.23 97.53 96.90 87.32 95.31 93.08 99.00 96.69 96.96 84.43 96.63 97.64 Deformable-DETR 96.68 98.42 97.98 75.17 95.53 95.58 99.00 97.55 96.95 91.91 96.62 98.04 Sparse-RCNN 96.38 98.37 97.69 62.11 96.22 95.86 99.01 97.79 97.84 88.39 97.29 97.97 TSRDet(Ours) 97.50 98.33 98.09 91.60 97.40 97.15 99.01 98.83 97.99 94.62 96.61 97.93 PubTables1M Cascade R-CNN 93.40 95.38 94.76 85.75 93.32 92.57 99.01 98.76 87.56 82.18 95.81 97.11 Deformable-DETR 94.82 97.43 96.79 78.33 92.55 94.48 98.99 97.89 95.84 85.04 95.43 95.74 Sparse-RCNN 96.46 98.14 97.60 84.25 95.73 96.45 99.00 98.42 98.03 87.85 97.91 97.57 TSRDet(Ours) 97.72 98.26 98.04 94.76 97.43 97.33 99.01 98.99 98.41 94.21 97.88 97.85
@@ 4.3 Ablation Study
In this section, we conduct experiments on the FinTabNet dataset to demonstrate the effectiveness of our applied methods, including applying aspect ratio tuning, pseudo-class generation, deformable convolution and spatial attention. The experimental results are shown in Tables 8 and 9, in which Asp_Ratio Tuning, Pseudo-Class Gen, DEFORM, and S_Attn are shorts for applying aspect ratio tuning, pseudo-class generation, deformable convolution, and spatial attention, respectively. We use the default implementation of Cascade R-CNN in Detectron2 wu2019detectron2 as the baseline model. Even though this default implementation can reach 95.06% regarding the mAP, its overall structure-only TEDS only reaches 82.70%. After tuning the aspect ratios for the anchor generation, the structure-only TEDS is increased to 90.23%, even though the mAP is only increased from 95.06% to 95.54%. Other ablation results show similar trends compared with the baseline model. It is worth mentioning that simply applying deformable convolution without other methods can increase the detection performance significantly, but lead to worse structure-only TEDS score, , which is a similar phenomenon aforementioned in Section 4.2. This is because deformable convolution can make model fit better to the extra spaces that are not necessarily useful for recovering the table structure.
Table 8: Ablation study results on FinTabNet dataset with structure-only TEDS score. Asp_Ratio Tuning, Pseudo-Class Gen, DEFORM, and S_Attn are shorts for applying aspect ratio tuning, pseudo-class generation, deformable convolution, and spatial attention.
Model Asp_Ratio Tuning Pseudo-Class Gen DEFORM S_Attn TEDS-struc.(%)
@@ 4.4 Observations and Analysis
As mentioned earlier in Sections 4.2 and  4.3, there are some important observations from the experimental results that can demonstrate the key design aspects included in this study. In this section, we further discuss and explain these important observations.
4.4.1 Metrics misalignment
The experimental results in Sections 4.2 and  4.3 show that a model with better mAP performance does not necessarily also have better structural TEDS, which means that the structure-only TEDS and the COCO metrics are not aligned. For example, in Table 7, both the Deformable-DETR and our proposed TSRDet can reach 96.28% regarding mAP on the SciTSR dataset, which is better than that of Sparse R-CNN. However, when it comes to structure-only TEDS, as shown in Table 3, both TSRDet and Sparse R-CNN can perform better than Deformable-DETR. COCO metrics evaluate the similarity between the prediction and ground truth boxes with IoU thresholds. In contrast, the structure-only TEDS evaluates the structure difference between the detected components and the ground truth components because the post-processing method uses structure information, such as the overlaps between columns and rows, to generate table cells. In the context of detection-based TSR, the defined components often contain large white areas that can influence the COCO metrics, but cannot influence the structure. Figure 8 shows the columns’ ground truth of a sample from the FinTabNet dataset. As shown in Figure 7(a), the minimum boxes required to infer the correct architecture are much smaller than the ground truth boxes. Figure 7(b) and  7(c) show two prediction samples after post-processing using Ablation 1 and Ablation3 models. Ablation 1 model with deformable convolution can better fit the extra white areas to improve the detection performance, but it is not necessarily helpful for the structure.
(a) Comparison of the ground truth bounding box and the minimum bounding box for the structure. We only include the columns’ boxes for simplicity.
(b) A sample prediction result of the Ablation 1 model after post-processing. The mAP and structure-only TEDS are 97.22 and 84.35, respectively. We only include the columns’ predictions for simplicity.
(c) A sample prediction result of the Ablation 3 model after post-processing. The mAP and structure-only TEDS are 95.51 and 96.95, respectively. We only include the columns’ predictions for simplicity.

### 10 Jan 2024 v2

@@ 4.3 Spatial Attention and Deformable Convolution
As discussed in section 3.5, building long-range dependencies for detecting the defined components is important. Inspired by the recent studies using large convolution kernels ding2022scaling ; liu2022convnet , we introduce a Spatial Attention Module for our solution, whose architecture is shown in Figure 8. For the design of the Spatial Attention Module, we use a similar architecture with MSCA guo2022segnext containing multiple branches and large kernel convolutions and use spatial and depthwise separable convolution chollet2017xception ; howard2017mobilenets to reduce the number of parameters. More specifically, for the spatial separable convolution, we use a pair of $7*1$ and $1*7$ kernels to replace a typical $7*7$, use the pair of $11*1$ and $1*1$ 11, and the pair of $21\times 1$ and $1\times 21$ to replace $11*1$ 11 and $21*21$ kernels, respectively. For the depthwise separable convolution, we applied the convolution on each channel of the feature maps independently. Then, the outputs of the three branches are concatenated together as the input of a convolution layer with $1\times 1$ kernel to make the channel dimension the same as the inputs. The proposed Spatial Attention Module can be easily inserted into the Backbone Network between two blocks because they do not change the feature shapes. For example, for a typical backbone network implemented by ResNet he2016deep containing a STEM block and four Residual Blocks, as shown in Figure 8, the Spatial Attention Module can be inserted after the last three Residual Blocks to generate the spatial attention, then the spatial attention can be applied to the original outputs of each Residual Block by Element-wise Multiplication. It is worth mentioning that the Spatial Attention Module shown in Figure 8 have independent trainable parameters, and all the feature maps are padded correspondingly to keep the size of the feature maps.
@@
On the other hand, as discussed in section 3.5, deformable convolution dai2017deformable can improve the local feature generation and has been demonstrated to help improve the detection performance on the document image detection tasks by many studies fernandes2023tablestrrec ; siddiqui2018decnt ; mondal2023dataset . Therefore, in this study, we apply the proposed Spatial Attention Module and deformable convolution to build long-range dependencies and improve local features together.
@@ 5.1 Datasets and Experimental Settings
We utilize four datasets in this work, including SciTSR chi2019complicated , FinTabNet zheng2020global , PubTabNet zhong2020image and PubTables1M smock2022pubtables . As discussed in the study smock2023aligning , FinTabNet and SciTSR datasets contain noise annotations that harm the model performance. Therefore, we use their cleaned versions proposed in the study in smock2023aligning . Each image sample in these four datasets contains only a table with extra padding pixels to ensure the entire table is extracted. The SicTSR dataset is collected from academic publications containing 7453, 1034, and 2134 samples for training, validation, and testing. PubTables1M dataset is a large-scale dataset for the TSR problem collected from the PMCOA corpus, containing 758849 training samples, 94959 validation samples, and 93834 testing samples. Since the PubTabNet dataset does not provide original PDF files, we cannot process it to make detection annotations. Besides, its testing is not publicly available. Therefore, we use its validation set to evaluate the model trained with the PubTable1M dataset. Following the study in smock2023aligning , we use the code base in smock2021tabletransformer to process the datasets and align the formats of these datasets. FinTabNet is also a large dataset widely used for the TSR problem, containing 78537, 9289, and 9650 samples for training, testing, and validation. FinTabNet is collected from the annual reports of companies, making its data source different from the other datasets. Table 1 summarizes the datasets used in this study for the model evaluation.
@@
Since the TSR problem in this study is formulated as an object detection problem, we use both detection and cell-level TSR metrics for the model evaluation. For the detection metric, we employ the widely accepted COCO metrics lin2014microsoft , which has been discussed in section 3.4. More specifically, mean Average Precision (AP), $AP50$, $AP75$, $AP_{s}$, $AP_{m}$, $AP_{l}$, and object-specific AP scores are used as metrics, where $AP50$, $AP75$ are the APs using 0.50 and 0.75 as IoU thresholds, respectively. $AP_{s}$, $AP_{m}$, and $AP_{l}$ are the APs of different target object sizes, defined by Equation 12.
@@
For the TSR metric, we choose structure-only Tree-Edit-Distance-Based Similarity(TEDS) zhong2020image , which is firstly introduced in the study to overcome the drawbacks of adjacency relation metrics hurst2003constraint , and can be defined as Equation 8 as discussed in section 3.4. We use structure-only TEDS in the study, which only considers the HTML tags without extracting their contents to avoid the influence of OCR tools. The testing samples can be categorized into simple and complex groups based on whether they have cells spanning multiple columns and rows.
@@ 5.2 Implementation Details and Experimental Results
To verify the effectiveness of our proposed solution, we include all three types of state-of-the-art methods as discussed in section 2. For the detection-based methods, Cascade R-CNN Cai2018CascadeRD , Deformable-DETR zhu2020deformable and Sparse R-CNN sun2021sparse are used as benchmark models, in which Cascade R-CNN Cai2018CascadeRD is also the based model of the proposed methods, Deformable-DETR and Sparse R-CNN are two state-of-the-art transformer-based detection models. For the image-to-sequence models, EDD zhong2020image , TableFormer nassar2022tableformer , TableMaster ye2021pingan , VAST huang2023improving and MTL-TabNet visapp23namly are included. TSRFormer-DQ-DETR WANG2023109817 and RobustTabNet ma2023robust are two state-of-the-art models following the pipeline of detecting separation lines and then merging cell grids, which can treated as a graph-based model as discussed in section 2. TSRNet li2022table is also a graph-based methods which detect table cells first, then applys GNN to build the relations among the detected cells.
@@
PRE_NMS_TOPK_TRAIN 4000 RPN proposals to keep before applying NMS in training
PRE_NMS_TOPK_TEST 2000 RPN proposals to keep before applying NMS in testing
POST_NMS_TOPK_TRAIN 4000 RPN proposals to keep after applying NMS in training
POST_NMS_TOPK_TEST 2000 RPN proposals to keep after applying NMS in testing
@@
We implement Cascade R-CNN and our proposed method based on the Detection2 wu2019detectron2 , the Deformable-DETR based on detrex ren2023detrex , and the Sparse R-CNN with their official codebase. For the Deformable-DETR and Sparse R-CNN, we use their default parameters. For the Cascade R-CNN baseline, we align the number of regional proposals and the batch normalization method to the TSRDet, as shown in Table 2. All these detection models are using ResNet50 he2016deep pre-trained with ImageNet deng2009imagenet as the backbone network. We also re-train TableMaster ye2021pingan with the FinTabNet dataset based on its official code base. We term our proposed method with TSRDet for fast reference. For the implementation of the proposed TSRDet, aspect ratios in the anchor box generation are set as [0.0125, 0.025, 0.0625, 0.125, 0.25, 0.5, 1.0, 2.0, 4.0, 8.0, 16, 40, 80], and other key parameters are summarized in Table 2. Notably, to calculate the structure-only TEDS, we use the scripts provided by study in smock2021tabletransformer to generate the HTML sequences from the detected components, and all benchmark models, except our proposed model, are using the original definition of PubTables1M, which treats all table components independently with its multi-label detection setting. All the models are trained with 240, 120, and 60 epochs for the SciTSR, FinTabNet, and PubTables1M datasets, respectively.
@@
The experimental results regarding the structure-only TEDS and COCO metrics are shown in Tables 3456 and 7, which can demonstrate the superiority of the proposed solution. For the SciTSR dataset, the proposed TSRDet can improve the baseline Cascade R-CNN by 19.32% regarding the structure-only TEDS, outperforming Deformable-DETR and Sparse R-CNN. When it comes to COCO metrics, the mAP of the proposed TSR is as good as Deformable-DETR, outperforming other benchmark models. Similarly, the proposed TSRDet can also outperform benchmark models regarding both COCO metrics and structure-only TEDS on the FinTabNet and PubTables1M datasets. It is worth mentioning that PubTabNet dataset does not provide original PDF files, making it hard to generate detection annotations. Therefore, the model performance reported in Table 6 is calculated using the model trained with the PubTable1M dataset. Even though the PubTable1M and PubTabNet datasets have misalignments regarding the ground truth HTML sequences, the proposed method still shows competitive performance compared with other state-of-the-art methods, as shown in Table 6. Figure 9 shows a prediction sample and its generated HTML sequence after post-processing, demonstrating the capacities of the proposed solution.
@@
Dataset Model $AP$ $AP50$ $AP75$ $AP_{s}$ $AP_{m}$ $AP_{l}$ Table Column Row Spanning Cell Projected Row Header Column Header SciTSR Cascade R-CNN 93.89 95.27 94.80 95.81 93.89 92.96 98.96 98.63 96.33 88.58 83.80 97.01 Deformable-DETR 96.28 97.39 97.01 96.75 96.55 96.07 98.96 98.63 97.26 93.84 90.86 98.15 Sparse R-CNN 94.78 96.17 95.48 95.49 95.07 90.08 98.98 98.30 97.93 88.06 86.92 98.49 TSRDet(Ours) 96.28 96.79 96.57 99.01 96.42 95.65 98.97 99.25 98.57 95.30 87.06 98.50 FinTabNet Cascade R-CNN 95.23 97.53 96.90 87.32 95.31 93.08 99.00 96.69 96.96 84.43 96.63 97.64 Deformable-DETR 96.68 98.42 97.98 75.17 95.53 95.58 99.00 97.55 96.95 91.91 96.62 98.04 Sparse R-CNN 96.38 98.37 97.69 62.11 96.22 95.86 99.01 97.79 97.84 88.39 97.29 97.97 TSRDet(Ours) 97.50 98.33 98.09 91.60 97.40 97.15 99.01 98.83 97.99 94.62 96.61 97.93 PubTables1M Cascade R-CNN 93.40 95.38 94.76 85.75 93.32 92.57 99.01 98.76 87.56 82.18 95.81 97.11 Deformable-DETR 94.82 97.43 96.79 78.33 92.55 94.48 98.99 97.89 95.84 85.04 95.43 95.74 Sparse R-CNN 96.46 98.14 97.60 84.25 95.73 96.45 99.00 98.42 98.03 87.85 97.91 97.57 TSRDet(Ours) 97.72 98.26 98.04 94.76 97.43 97.33 99.01 98.99 98.41 94.21 97.88 97.85
@@ 5.3 Ablation Study
In this section, we conduct experiments on the FinTabNet dataset to demonstrate the effectiveness of our applied methods, including using the proposed single-label detection formulation, tuning parameters of RPN, applying the deformable convolution and spatial attention. It is worth mentioning that tuning parameters of RPN includes increasing the number of proposals and ajusting the aspect rations. Since other studies have successfully applied the effectiveness of increasing the number of proposals, we applied it to both the Cascade R-CNN baseline and the proposed TSRDet, as discussed in section 3.3 and 5.2. Therefore, we only discuss the the impact of adjusting the aspect ratios for tuning parameters of RPN in this section.
The experimental results are shown in Tables 8 and 9, in which Asp_Ratio Tuning, Single_Label, DEFORM, and S_Attn are shorts for applying aspect ratio tuning, single label formulation, deformable convolution, and spatial attention, respectively. Even though Cascade R-CNN baseline can reach 95.06% regarding the mAP, its overall structure-only TEDS only reaches 82.70%. After tuning the aspect ratios for the anchor generation, the structure-only TEDS is increased to 90.23%, even though the mAP is only increased from 95.06% to 95.54%. Applying deformable convolution without other methods can improve the detection performance significantly but lead to a worse structure-only TEDS if we compare Ablation 1 and the Cascade R-CNN baseline. Ablation 3 and Ablation 4 show that transforming the multi-label detection formulation into single-label formulation can significantly improve the performance, and also make deformable convolution improve the model performance. And applying both deformable convolution and spatial attention together can further improve the model performance from the results of Abliation 4, 5 and TSRDet, as shown in Table 8. On the other hand, when it comes to detection metrics, applying deformable convolution always brings performance improvements from the results of Ablation 1 and Ablation 4, which can verify our analysis on the mismatch of detection metrics and cell-level metrics in section 3.4.
Table 8: Ablation study results on FinTabNet dataset with structure-only TEDS score. Asp_Ratio Tuning, Single_Label, DEFORM, and S_Attn are shorts for applying aspect ratio tuning, single-label formulation, deformable convolution, and spatial attention.
Model Asp_Ratio Tuning Single_Label DEFORM S_Attn TEDS-struc.(%)
@@ 5.4 Observations and Analysis
Sections 5.2 and 5.3 have demonstrated the effectiveness of our proposed methods. In this section, we further discuss some observations from the experimental results and how these observations verify our analysis in section 3.
5.4.1 Multi-label Detection
As discussed in sections 3.1 and 3.2, multi-label detection tasks are difficult for two-stage detection models, but transformer-based detection models with learnable proposals can deal with multi-label detection tasks. Besides, the problem formulation of PubTables1M is a multi-label task, making it difficult for two-stage detection models. The experimental results from sections 5.2 can demonstrate our analysis. For example, as shown in Table 3, the performance of Deformable-DETR and Sparse R-CNN are 97.81% and 98.13% regarding the structure-only TEDS, which are very close to the performance of proposed TSRDet (98.41%) and far better than the Cascade R-CNN baseline (79.09%). Notably, as mentioned in section 5, all the models, except our proposed TSRDet, are using the multi-label detection setting. Therefore, two trasnformer-based detection models show promising results in the multi-label detection setting. Similarly, the experiments on the FinTabNet and PubTables1M datasets also show similar results. For example, the structure-only TEDS performance of Sparse R-CNN, TSRDet, and Cascade R-CNN baseline are 97.72%, 98.55%, and 83.78% on the PubTables1M dataset, 98.13, 99.05, and 87.49 on the FinTabNet dataset.
5.4.2 The Misalignment of Metrics
The experimental results in sections 5.2 and 5.3 show the misalignment of COCO and TEDS metrics many times. For example, in Table 7, both the Deformable-DETR and our proposed TSRDet can reach 96.28% regarding mAP on the SciTSR dataset, which is better than that of Sparse R-CNN. However, when it comes to structure-only TEDS, as shown in Table 3, both TSRDet and Sparse R-CNN can perform better than Deformable-DETR. Similar results also appear in the experiments on the FinTabNet dataset. As shown in Table 7, the mAP of Sparse-RCNN and Deformable-DETR are 96.38% and 96.68%, while their structure-only TEDS are 97.81% and 98.13%. More similar results can be found in the results of the ablation study, such as Ablation 1 and Ablation 3, as shown in Tables 8 and 9. To further verify our discussion in section 3.4, we show the prediction results of Ablation 1 and 3 in Figure 10. As discussed in section 3.4, COCO metrics are relied on IoU scores, while TEDS is not. Therefore, as shown in Figure 9(a), Ablation 1 with deformable convolution can better fit the extra white areas to improve the detection performance, but it cannot improve the TEDS compared with Ablation 3 whose result is shown in Figure 9(b). It is worth mentioning that the ground truth of the sample in Figure 10 has been shown in Figure 5.
(a) A sample prediction result of the Ablation 1 model. The mAP and structure-only TEDS are 97.22 and 84.35, respectively. We only include the columns’ predictions for simplicity.
(b) A sample prediction result of the Ablation 3 model. The mAP and structure-only TEDS are 95.51 and 96.95, respectively. We only include the columns’ predictions for simplicity.